OpenStudy (anonymous):

Can someone PLEASE help me with 2 log problems. i really need help

OpenStudy (unklerhaukus):

yeah, where are the problems

OpenStudy (unklerhaukus):

Here are some properties of logarithms: \[\log_bx=y\quad\iff\quad x=b^y\] \[\log_b(x)+\log_b(y)=\log_b(xy)\]\[\log_b(x)-\log_b(y)=\log_b(x/y)\] \[\log_b(b)=1\] \[\log_b(z^n)=n\log_b(z)\] \[\log_ba=\frac{\log_ca}{\log_cb}\]

OpenStudy (unklerhaukus):

@brittlynn, post you problems, then maybe some can help you

OpenStudy (anonymous):

heres one

OpenStudy (anonymous):

\[\log_{4}(\log_{2}16) \]

OpenStudy (anonymous):

theres the second one